OpenStudy (anonymous):

[9.01] The graph of which equation is shown below? a u-shaped graph on a coordinate plane that is opening up and has a vertex of (2, −3) y = x2 − 4x + 1 y = x2 + 4x + 1 y = −x2 − 4x + 1 y = −x2 + 4x + 1

OpenStudy (anonymous):

this is the graph

OpenStudy (anonymous):

help me please

OpenStudy (anonymous):

it is the graph of y = x2 − 4x + 1 y+3=x2 − 4x + 1+3 or y+3=(x-2)^2

OpenStudy (anonymous):

wht?

OpenStudy (anonymous):

please help me please

OpenStudy (anonymous):

@Hero @phi @agent0smith can someone please help me w/this

terenzreignz (terenzreignz):

A month-old question? D:

OpenStudy (anonymous):

Yea I have the same question on an online course

terenzreignz (terenzreignz):

Which isn't even yours? :3 anyway, check the vertices of each of the answer choices... If you have an equation of the form \[\Large y =\color{blue}ax^2 + \color{red}bx + \color{green}c\] Then the vertex of its graph is given by \[\huge \left(-\frac{\color{red}b}{2\color{blue}a}\qquad,\qquad\frac{4\color{blue}a\color{green}c-\color{red}b^2}{4\color{blue}a}\right)\]

terenzreignz (terenzreignz):

But it's opening upward, so that \(\large x^2\) coefficient MUST be positive, so that rules out choices C and D (since they both have \(\large -x^2\))

OpenStudy (anonymous):

so it's B because the 4 is positive?

terenzreignz (terenzreignz):

I don't actually know... so you know choice B opens upward, that's good. But what's the vertex of choice B? Use this formula: \[\huge \left(-\frac{\color{red}b}{2\color{blue}a}\qquad,\qquad\frac{4\color{blue}a\color{green}c-\color{red}b^2}{4\color{blue}a}\right)\]

terenzreignz (terenzreignz):

\[\Large y = x^2 + 4x + 1 = \color{blue}1x^2 + \color{red}4x+\color{green}{1}\]

OpenStudy (anonymous):

so x=-2

terenzreignz (terenzreignz):

Yup... but you want the vertex to be (2,-3) so...?

OpenStudy (anonymous):

and the second equation is -4 I think

terenzreignz (terenzreignz):

It isn't, but anyway, the first coordinate, -b/2a gives you -2. But the vertex that you want is (2,-3)

OpenStudy (anonymous):

y=-3

OpenStudy (anonymous):

how do I change x to positive

terenzreignz (terenzreignz):

LOL You don't ^_^ You just figured out (though you seem to not know) the vertex of the choice B Which is x = -2, y = -3 Or (-2,-3) But that's something you can't change... the vertex of the second equation IS (-2,-3) which means it isn't the equation you're looking for :P

OpenStudy (anonymous):

So it's A

terenzreignz (terenzreignz):

By process of elimination, it must be :) Though you'd do well to apply the vertex formula to choice A \[\huge \left(-\frac{\color{red}b}{2\color{blue}a}\qquad,\qquad\frac{4\color{blue}a\color{green}c-\color{red}b^2}{4\color{blue}a}\right)\] Just to assure yourself that it IS the correct answer :)

terenzreignz (terenzreignz):

And that you SHOULD get the vertex x = 2 and y = -3 (2 , -3) Cheers ^_^

OpenStudy (anonymous):

Thanks so much

terenzreignz (terenzreignz):

Don't forget why choices C and D were not considered, okay? The question says the graph opens upward, and this is determined solely by the sign of the \(\large x^2\) term... if it's positive, as in choices A and B, the graph opens up, and if it's negative, as in choices C and D, the graph opens down ^_^